# More Than Memory Savings: Zeroth-Order Optimization Mitigates Forgetting in Continual Learning

### v1

@@ Abstract
Zeroth-order (ZO) optimization has gained attention as a memory-efficient alternative to first-order (FO) methods, particularly in settings where gradient computation is expensive or even impractical. Beyond its memory efficiency, in this work, we investigate ZO optimization for continual learning (CL) as a novel approach to address the plasticity-stability-efficiency trilemma. Through theoretical analysis and empirical evidence, we show that ZO optimization naturally leads to flatter loss landscapes, which in turn reduce forgetting in CL. However, this stability comes at a cost of plasticity: due to its imprecise gradient estimates and slower convergence, ZO optimization tends to be less effective than FO in acquiring new task-specific knowledge, particularly under constrained training budgets. To better understand this trade-off, we conduct a holistic evaluation of ZO optimization applied to various existing CL methods. Our findings reveal that ZO optimization enhances stability but often undermines plasticity, particularly when used with learnable classifiers. Motivated by this insight, we propose ZO-FC, a simple but effective approach that applies ZO optimization to a single adapter-based PEFT module with FO optimized classifier. This design leverages the stability benefits of ZO while preserving the adaptability of FO updates with negligible memory overhead. Experiments demonstrate that ZO-FC achieves an effective balance between stability and plasticity, offering a practical and memory-efficient solution for on-device CL.
22footnotetext: Corresponding author
@@ 1 Introduction
Emulating the human ability to learn continuously from new experiences is a key inspiration in artificial intelligence. This pursuit has driven significant attention to Continual Learning (CL) [17, 7, 47], a paradigm aiming to train models on a sequence of tasks while preserving previously acquired knowledge. A primary challenge of CL is to maintain a balance between plasticity, the ability to learn new knowledge, and stability, the ability to retain previously learned knowledge by preventing catastrophic forgetting: a phenomenon where models lose prior knowledge when training on new tasks [13].
@@
To address this plasticity-stability dilemma, recent progress in CL has shifted from “continual training from scratch” to a new paradigm of “continual fine-tuning from pre-trained models”. In particular, driven by the increasingly powerful pre-trained foundation models, such as Vision Transformers (ViTs) [8, 43], and the effectiveness of parameter-efficient fine-tuning (PEFT) techniques (i.e., prompts [24], adapters [19], or LoRAs [20]), recent CL methods [48, 15, 53, 54, 25] follow the recipe to freeze the backbone and train lightweight modules across tasks. This setup inherits strong plasticity that new tasks are learned quickly from a good initialization so the critical challenge becomes mitigating catastrophic forgetting to enhance stability. These methods therefore isolate task-specific information into PEFT modules to minimize interference with previously learned tasks.
@@
Recently, zeroth-order (ZO) optimization has shown promising results in fine-tuning large language models [29, 52], providing a memory-efficient alternative to traditional first-order (FO) optimization while achieving comparable accuracy. ZO optimization eliminates the need for backpropagation and the associated storage of intermediate activations, making training memory close to inference. In this work, we target low-memory, on-device CL settings and further look beyond this established memory advantage. We hypothesize that ZO’s gradient-free search makes it naturally tend to wider, flatter minima, which is known to be more robust to catastrophic forgetting [12, 38]. Motivated by this potential, we investigate ZO optimization for CL as a novel approach to address the plasticity-stability-efficiency trilemma, aiming to achieve accuracy comparable to FO methods at a fraction of the training memory overhead. To this end, our contribution can be summarized as follows:
@@ 2.1 PEFT-based Class-incremental Learning
Class-incremental learning (CIL) is a widely used and challenging setting in continual learning, where new groups of classes are introduced sequentially, and the model must learn to classify all previously seen classes without access to task or class identity during inference [55]. With the growing adoption of large-scale pre-trained foundation models (e.g., Vision Transformer [8, 43]), a common learning paradigm in CIL is to use the pre-trained model as a strong initialization and perform continual fine-tuning on sequential tasks using parameter-efficient fine-tuning (PEFT), while keeping the pre-trained backbone frozen. PEFT enables efficient adaptation by introducing lightweight trainable modules, with common techniques including prompts [24], adapters [19], and LoRAs [20]. These PEFT-based CIL methods [48, 15, 53, 54, 25] have shown strong performance, even under rehearsal-free settings where no data from previous tasks is used during the training of new tasks.
By leveraging strong pre-trained models as initialization, PEFT-based CIL methods inherently exhibit high plasticity. As a result, their primary focus is on mitigating forgetting to enhance stability in learning each task. First of all, SimpleCIL [53] achieves good accuracy, particularly on simpler datasets such as CIFAR100, by only applying prototype-based classification on top of a frozen pre-trained model without any learning. This result empirically highlights the effectiveness of leveraging pre-trained representations. Moreover, to mitigate forgetting when introducing task adaptive PEFT modules, a series of methods has been proposed. L2P [48] uses a learnable pool of prompts to activate task-specific representations, effectively isolating knowledge across tasks. CODA-Prompt [40] refines L2P by applying contrastive alignment to ensure consistent representation. LAE [15] maintains an online PEFT module for current-task learning and an offline PEFT module updated via exponential moving average (EMA) so that these modules can be ensembled at inference to combine plastic adaptation and stable accumulation. APER [53] extends SimpleCIL to adapt the pre-trained model only during the first task, then merges features from both the frozen pre-trained backbone and the adapted model to form a merged feature extractor. EASE [54] adopts a task-specific adapter for each task, and introduces a semantic-guided prototype synthesis strategy that reconstructs old class representations in the current subspace to mitigate forgetting. InfLoRA [25] adapts LoRA-based CL with an interference-free design by constraining the update directions of LoRA modules to be task-orthogonal to reduce cross-task interference. In this work, these methods serve as FO optimization baselines for evaluating the application of ZO optimization.
@@ 2.2 Zeroth-Order Optimization
Zeroth-order (ZO) optimization has been extensively studied [41, 32, 26, 28] as a gradient-free alternative to first-order (FO) methods. By estimating gradients using only function evaluations without the need for backpropagation, ZO optimization is appealing in privacy-sensitive, black-box, or resource-constrained settings, such as federated learning and edge device deployment [4, 45, 10, 39]. A representative ZO method is Simultaneous Perturbation Stochastic Approximation (SPSA) [41], which estimates gradients by computing finite differences of the loss function along randomly sampled direction vectors. Given model parameters $\theta$ and a loss function $L$, the ZO gradient estimate can be expressed as:
@@
where $\Delta\sim{\pm 1}^{d}$ is a random perturbation vector, $d$ is the parameter dimension, and $\varepsilon$ is a small perturbation scale. This ZO estimate can be further incorporated into standard optimizers, such as stochastic gradient descent (SGD), to yield batch-wise zeroth-order stochastic gradient descent (ZO-SGD) [16, 32]. To reduce the high variance inherent in single-direction estimates for each updating step, a common practical extension is to average over $Q$ queries of independent perturbation directions per gradient step [42].
Notably, recent work MeZO [29] has demonstrated that ZO-SGD fine-tuning on large language models can achieve accuracy comparable to its FO counterpart. In addition, while there are other ZO techniques, including Forward-Grad [1, 36], Sign-based variants [27, 29], and adaptive methods such as ZO-AdaMM [6], recent benchmark studies [52, 11] show that no single method consistently outperforms others in all tasks. Therefore, we adopt SPSA and its mini-batch extension, ZO-SGD, as our default ZO optimization strategy.
@@ 3 Rethinking Zeroth-Order Optimization for Continual Learning
1) Loss Flatness Mitigates Forgetting: Let $\theta_{t}\!\in\!\mathbb{R}^{d}$ denote the model parameters after training on task $t$. When learning task t+1t{+}1, the parameters are updated by $\Delta\theta=\theta_{t+1}-\theta_{t}$. Following [31], we quantify forgetting on previous tasks as the increase in their cumulative loss:
@@
where $H_{t}=\nabla^{2}L_{\text{old}}(\theta_{t})$, representing the Hessian of the old-task loss. As proved by prior works [30], the forgetting can be bounded by:
@@
where $\lambda_{\max}(H_{t})$ is the largest eigenvalue of the Hessian matrix at task $t$. As revealed by [21], this eigenvalue captures the sharpness of the loss landscape: a higher $\lambda_{\max}(H_{t})$ indicates a sharper curvature, meaning that small changes in parameters $\Delta\theta$ can result in large increases in loss. As a result, models operating in sharper regions are more prone to catastrophic forgetting, since even minor updates can severely disrupt performance on previously learned tasks.
Furthermore, recent work [22] empirically confirms that smaller curvature (i.e., lower $\lambda_{\max}(H_{t})$) corresponds to flatter regions of the loss surface, which are associated with reduced forgetting in continual learning. Motivated by this, several studies [31, 49, 2, 38] have proposed augmenting standard FO optimization with auxiliary loss terms or steps aimed at minimizing curvature surrogates to promote flatter minima. In contrast, we aim to explore mitigating forgetting through the intrinsic bias of ZO optimization.
2) ZO optimization tends to flatter minima: Foundational work [32] has theoretically shown that every ZO optimization step (Eq.1) is, in fact, an update on a smoothed version of the original loss. For any twice-differentiable loss $L(\theta)$ and an isotropic random direction $\Delta$, the smooth loss is defined as the following expected loss:
@@
where the noise term $O(\varepsilon^{4})$ is negligible. The extra term $\frac{\varepsilon^{2}}{2}\operatorname{Tr}(\nabla^{2}L)$ is proportional to the trace of the Hessian, which can be seen as a measure of curvature of the loss landscape [51]. Minimizing $L_{\varepsilon}$ through ZO optimization therefore minimizes the original loss and simultaneously penalizes large positive curvature. Because sharp minima possess large positive Hessian trace, they receive a larger penalty in $L_{\varepsilon}$. ZO SPSA updates, which descend on this smoothed objective, are statistically steered away from narrow valleys and into flatter regions of the landscape. When ZO optimization converges to a stationary point of $L_{\varepsilon}$, the trace of the Hessian of the original loss at that point is also minimized. Therefore, ZO optimization inherently favors flat minima without the need for explicit regularization.
@@
However, this benefit comes with a trade-off in plasticity. Unlike FO optimization, which efficiently converges toward local minima via accurate gradient information, ZO optimization relies on stochastic perturbations and typically exhibits slower convergence. Specifically, the convergence rate of ZO methods is $O(d/Q)$, worse than their FO counterparts [16], where $d$ is the parameter dimension and $Q$ is the number of random queries per step, making them less efficient than FO [32]. As a result, given the limited training budget for CL, this slow convergence may prevent ZO methods from reaching sufficiently good local minima, thus impairing plasticity and overall task adaptation.
@@ 4.1 Direct ZO Substitution in CL Fails
Prototype[34] (nearest-centroid) Centroid: $c_{y}=\dfrac{1}{|D_{y}|}\sum\!f_{\theta}(x)$ Predict: $\hat{y}=\arg\min_{y}\|h-c_{y}\|^{2}$ SimpleCIL[53] EASE[54] APER_Adapter[53]
Unified learnable (linear / cosine) Linear: $s_{y}(h)=w_{y}^{\!\top}h$ Cosine[46]: $s_{y}(h)=\alpha\dfrac{w_{y}^{\!\top}h}{\|w_{y}\|\|h\|}$ Predict: $\hat{y}=\arg\max_{y}s_{y}(h)$ Adapter[5] LAE[15] L2P[48] CODA-Prompt[40] InfLoRA[25]
@@ 4.2 ZO-FC: ZO optimization for PEFT module with FO learnable classifier
Although ZO-FC employs FO optimization for the classifier, it still benefits from the flat minima effect: the ZO updates bias the adapter representation toward flatter solutions, while the FO classifier preserves decision-boundary plasticity (proof in Appendix C). Besides, its impact on memory efficiency remains negligible. This is because the classifier is the final layer in the model and does not propagate gradients back to earlier layers. As a result, intermediate activations from the backbone and adapter do not need to be stored during training, significantly reducing memory usage. Therefore, ZO-FC retains the key advantage of memory efficiency while benefiting from the precision of first-order updates where they matter most. The full procedure is detailed in Algorithm 1 in Appendix B.

### v2

@@ Abstract
Zeroth-order (ZO) optimization has gained attention as a memory-efficient alternative to first-order (FO) methods, particularly in settings where gradient computation is expensive or even impractical. Beyond its memory efficiency, in this work, we investigate ZO optimization for continual learning (CL) as a novel approach to address the plasticity-stability-efficiency trilemma. Through theoretical analysis and empirical evidence, we show that ZO optimization naturally leads to flatter loss landscapes, which in turn reduce forgetting in CL. However, this stability comes at a cost of plasticity: due to its imprecise gradient estimates and slower convergence, ZO optimization tends to be less effective than FO in acquiring new task-specific knowledge, particularly under constrained training budgets. To better understand this trade-off, we conduct a holistic evaluation of ZO optimization applied to various existing CL methods. Our findings reveal that ZO optimization enhances stability but often undermines plasticity, particularly when used with learnable classifiers. Motivated by this insight, we propose ZO-FC, a simple but effective approach that applies ZO optimization to a single adapter-based PEFT module with FO optimized classifier. This design leverages the stability benefits of ZO while preserving the adaptability of FO updates with negligible memory overhead. Experiments demonstrate that ZO-FC achieves an effective balance between stability and plasticity, offering a practical and memory-efficient solution for on-device CL. Our code is available at: https://github.com/Brian-WY/ZOFC.
footnotetext: Li Yang is the corresponding author.
@@ 1 Introduction
Emulating the human ability to learn continuously from new experiences is a key inspiration in artificial intelligence. This pursuit has driven significant attention to Continual Learning (CL) [hadsell2020embracing, de2021continual, wang2024comprehensive], a paradigm aiming to train models on a sequence of tasks while preserving previously acquired knowledge. A primary challenge of CL is to maintain a balance between plasticity, the ability to learn new knowledge, and stability, the ability to retain previously learned knowledge by preventing catastrophic forgetting: a phenomenon where models lose prior knowledge when training on new tasks [FRENCH1999128].
@@
To address this plasticity-stability dilemma, recent progress in CL has shifted from “continual training from scratch” to a new paradigm of “continual fine-tuning from pre-trained models”. In particular, driven by the increasingly powerful pre-trained foundation models, such as Vision Transformers (ViTs) [dosovitskiy2021an, steiner2021train], and the effectiveness of parameter-efficient fine-tuning (PEFT) techniques (i.e., prompts [lester2021prompt], adapters [houlsby2019adapter], or LoRAs [hu2022lora]), recent CL methods [wang2022l2p, gao2023lae, zhou2024aper, zhou2024ease, liang2024inflora] follow the recipe to freeze the backbone and train lightweight modules across tasks. This setup inherits strong plasticity that new tasks are learned quickly from a good initialization so the critical challenge becomes mitigating catastrophic forgetting to enhance stability. These methods therefore isolate task-specific information into PEFT modules to minimize interference with previously learned tasks.
@@
Recently, zeroth-order (ZO) optimization has shown promising results in fine-tuning large language models [malladi2023fine, zhang2024revisiting], providing a memory-efficient alternative to traditional first-order (FO) optimization while achieving comparable accuracy. ZO optimization eliminates the need for backpropagation and the associated storage of intermediate activations, making training memory close to inference. In this work, we target low-memory, on-device CL settings and further look beyond this established memory advantage. We hypothesize that ZO’s gradient-free search makes it naturally tend to wider, flatter minima, which is known to be more robust to catastrophic forgetting [foret2021sharpnessaware, shi2021fclflat]. Motivated by this potential, we investigate ZO optimization for CL as a novel approach to address the plasticity-stability-efficiency trilemma, aiming to achieve accuracy comparable to FO methods at a fraction of the training memory overhead. To this end, our contribution can be summarized as follows:
@@ 2.1 PEFT-based Class-incremental Learning
Class-incremental learning (CIL) is a widely used and challenging setting in continual learning, where new groups of classes are introduced sequentially, and the model must learn to classify all previously seen classes without access to task or class identity during inference [zhu2021class]. With the growing adoption of large-scale pre-trained foundation models (e.g., Vision Transformer [dosovitskiy2021an, steiner2021train]), a common learning paradigm in CIL is to use the pre-trained model as a strong initialization and perform continual fine-tuning on sequential tasks using parameter-efficient fine-tuning (PEFT), while keeping the pre-trained backbone frozen. PEFT enables efficient adaptation by introducing lightweight trainable modules, with common techniques including prompts [lester2021prompt], adapters [houlsby2019adapter], and LoRAs [hu2022lora]. These PEFT-based CIL methods [wang2022l2p, gao2023lae, zhou2024aper, zhou2024ease, liang2024inflora] have shown strong performance, even under rehearsal-free settings where no data from previous tasks is used during the training of new tasks.
By leveraging strong pre-trained models as initialization, PEFT-based CIL methods inherently exhibit high plasticity. As a result, their primary focus is on mitigating forgetting to enhance stability in learning each task. First of all, SimpleCIL [zhou2024aper] achieves good accuracy, particularly on simpler datasets such as CIFAR100, by only applying prototype-based classification on top of a frozen pre-trained model without any learning. This result empirically highlights the effectiveness of leveraging pre-trained representations. Moreover, to mitigate forgetting when introducing task adaptive PEFT modules, a series of methods has been proposed. L2P [wang2022l2p] uses a learnable pool of prompts to activate task-specific representations, effectively isolating knowledge across tasks. CODA-Prompt [smith2023codap] refines L2P by applying contrastive alignment to ensure consistent representation. LAE [gao2023lae] maintains an online PEFT module for current-task learning and an offline PEFT module updated via exponential moving average (EMA) so that these modules can be ensembled at inference to combine plastic adaptation and stable accumulation. APER [zhou2024aper] extends SimpleCIL to adapt the pre-trained model only during the first task, then merges features from both the frozen pre-trained backbone and the adapted model to form a merged feature extractor. EASE [zhou2024ease] adopts a task-specific adapter for each task, and introduces a semantic-guided prototype synthesis strategy that reconstructs old class representations in the current subspace to mitigate forgetting. InfLoRA [liang2024inflora] adapts LoRA-based CL with an interference-free design by constraining the update directions of LoRA modules to be task-orthogonal to reduce cross-task interference. In this work, these methods serve as FO optimization baselines for evaluating the application of ZO optimization.
@@ 2.2 Zeroth-Order Optimization
Zeroth-order (ZO) optimization has been extensively studied [spall1987spsa, nesterov2017gradientfree, liu2018zeroth, liu2020primer] as a gradient-free alternative to first-order (FO) methods. By estimating gradients using only function evaluations without the need for backpropagation, ZO optimization is appealing in privacy-sensitive, black-box, or resource-constrained settings, such as federated learning and edge device deployment [chen2017zoo, tu2019autozoom, fang2022communication, shu2023federated]. A representative ZO method is Simultaneous Perturbation Stochastic Approximation (SPSA) [spall1987spsa], which estimates gradients by computing finite differences of the loss function along randomly sampled direction vectors. Given model parameters $\theta$ and a loss function $L$, the ZO gradient estimate can be expressed as:
@@
where $\Delta\sim{\pm 1}^{d}$ is a random perturbation vector, $d$ is the parameter dimension, and $\varepsilon$ is a small perturbation scale. This ZO estimate can be further incorporated into standard optimizers, such as stochastic gradient descent (SGD), to yield batch-wise zeroth-order stochastic gradient descent (ZO-SGD) [ghadimi2013stochastic, nesterov2017gradientfree]. To reduce the high variance inherent in single-direction estimates for each updating step, a common practical extension is to average over $Q$ queries of independent perturbation directions per gradient step [spall2000adaptive].
Notably, recent work MeZO [malladi2023fine] has demonstrated that ZO-SGD fine-tuning on large language models can achieve accuracy comparable to its FO counterpart. In addition, while there are other ZO techniques, including Forward-Grad [baydin2022gradients, ren2022scaling], Sign-based variants [liu2019signsgd, malladi2023fine], and adaptive methods such as ZO-AdaMM [chen2019zo], recent benchmark studies [zhang2024revisiting, feng2025zeroflow] show that no single method consistently outperforms others in all tasks. Therefore, we adopt SPSA and its mini-batch extension, ZO-SGD, as our default ZO optimization strategy.
@@ 3 Rethinking Zeroth-Order Optimization for Continual Learning
1) Loss Flatness Mitigates Forgetting: Let $\theta_{t}\!\in\!\mathbb{R}^{d}$ denote the model parameters after training on task $t$. When learning task t+1t{+}1, the parameters are updated by $\Delta\theta=\theta_{t+1}-\theta_{t}$. Following [mirzadeh2021linear], we quantify forgetting on previous tasks as the increase in their cumulative loss:
@@
where $H_{t}=\nabla^{2}L_{\text{old}}(\theta_{t})$, representing the Hessian of the old-task loss. As proved by prior works [mirzadeh2020understanding], the forgetting can be bounded by:
@@
where $\lambda_{\max}(H_{t})$ is the largest eigenvalue of the Hessian matrix at task $t$. As revealed by [keskar2017largebatch], this eigenvalue captures the sharpness of the loss landscape: a higher $\lambda_{\max}(H_{t})$ indicates a sharper curvature, meaning that small changes in parameters $\Delta\theta$ can result in large increases in loss. As a result, models operating in sharper regions are more prone to catastrophic forgetting, since even minor updates can severely disrupt performance on previously learned tasks.
Furthermore, recent work [kong2024hessian] empirically confirms that smaller curvature (i.e., lower $\lambda_{\max}(H_{t})$) corresponds to flatter regions of the loss surface, which are associated with reduced forgetting in continual learning. Motivated by this, several studies [mirzadeh2021linear, wu2024cr, bian2024make, shi2021fclflat] have proposed augmenting standard FO optimization with auxiliary loss terms or steps aimed at minimizing curvature surrogates to promote flatter minima. In contrast, we aim to explore mitigating forgetting through the intrinsic bias of ZO optimization.
2) ZO optimization tends to flatter minima: Foundational work [nesterov2017gradientfree] has theoretically shown that every ZO optimization step (Eq.1) is, in fact, an update on a smoothed version of the original loss. For any twice-differentiable loss $L(\theta)$ and an isotropic random direction $\Delta$, the smooth loss is defined as the following expected loss:
@@
where the noise term $O(\varepsilon^{4})$ is negligible. The extra term $\frac{\varepsilon^{2}}{2}\operatorname{Tr}(\nabla^{2}L)$ is proportional to the trace of the Hessian, which can be seen as a measure of curvature of the loss landscape [zhang2025zofindsflat]. Minimizing $L_{\varepsilon}$ through ZO optimization therefore minimizes the original loss and simultaneously penalizes large positive curvature. Because sharp minima possess large positive Hessian trace, they receive a larger penalty in $L_{\varepsilon}$. ZO SPSA updates, which descend on this smoothed objective, are statistically steered away from narrow valleys and into flatter regions of the landscape. When ZO optimization converges to a stationary point of $L_{\varepsilon}$, the trace of the Hessian of the original loss at that point is also minimized. Therefore, ZO optimization inherently favors flat minima without the need for explicit regularization.
@@
However, this benefit comes with a trade-off in plasticity. Unlike FO optimization, which efficiently converges toward local minima via accurate gradient information, ZO optimization relies on stochastic perturbations and typically exhibits slower convergence. Specifically, the convergence rate of ZO methods is $O(d/Q)$, worse than their FO counterparts [ghadimi2013stochastic], where $d$ is the parameter dimension and $Q$ is the number of random queries per step, making them less efficient than FO [nesterov2017gradientfree]. As a result, given the limited training budget for CL, this slow convergence may prevent ZO methods from reaching sufficiently good local minima, thus impairing plasticity and overall task adaptation.
@@ 4.1 Direct ZO Substitution in CL Fails
Prototype[Rebuffi2016iCaRL] (nearest-centroid) Centroid: $c_{y}=\dfrac{1}{|D_{y}|}\sum\!f_{\theta}(x)$ Predict: $\hat{y}=\arg\min_{y}\|h-c_{y}\|^{2}$ SimpleCIL[zhou2024aper] EASE[zhou2024ease] APER_Adapter[zhou2024aper]
Unified learnable (linear / cosine) Linear: $s_{y}(h)=w_{y}^{\!\top}h$ Cosine[wang2018cosface]: $s_{y}(h)=\alpha\dfrac{w_{y}^{\!\top}h}{\|w_{y}\|\|h\|}$ Predict: $\hat{y}=\arg\max_{y}s_{y}(h)$ Adapter[chen2022adapter] LAE[gao2023lae] L2P[wang2022l2p] CODA-Prompt[smith2023codap] InfLoRA[liang2024inflora]
@@ 4.2 ZO-FC: ZO optimization for PEFT module with FO learnable classifier
Although ZO-FC employs FO optimization for the classifier, it still benefits from the flat minima effect: the ZO updates bias the adapter representation toward flatter solutions, while the FO classifier preserves decision-boundary plasticity (proof in supplementary material). Besides, its impact on memory efficiency remains negligible. This is because the classifier is the final layer in the model and does not propagate gradients back to earlier layers. As a result, intermediate activations from the backbone and adapter do not need to be stored during training, significantly reducing memory usage. Therefore, ZO-FC retains the key advantage of memory efficiency while benefiting from the precision of first-order updates where they matter most. The full procedure is detailed in Algorithm 1 in supplementary material.